# The Calibr-A-Ton: a novel method for calorimeter energy calibration

### v1

@@ Abstract
The energy calibration of calorimeters at collider experiments, such as the ones at the CERN Large Hadron Collider, is crucial for achieving the experiment’s physics objectives. Standard calibration approaches have limitations which become more pronounced as detector granularity increases. In this paper we propose a novel calibration procedure to simultaneously calibrate individual detector cells belonging to a particle shower, by targeting a well-controlled energy reference. The method bypasses some of the difficulties that exist in more standard approaches. It is implemented using differentiable programming. In this paper, simulated energy deposits in the electromagnetic section of a high-granularity calorimeter are used to study the method and demonstrate its performance. It is shown that the method is able to correct for biases in the energy response.
@@ 1 Introduction
Energy calibration of a calorimeter consists in methods that allow to precisely estimate the energy of particles that deposit energy within its volume. An imprecise calorimeter calibration can induce biases and uncertainties on the measurement of particles energies and hence, can significantly reduce the discovery potential of a physics experiment.
@@
Since calorimeters are segmented into detection sub-components, commonly called “cells”, the first step of an energy calibration consists in finding, for each cell, the factor that allows a precise correspondence between the signal given by that cell and the genuine energy deposit that occurred in that cell. Those factors are typically called “calibration constants”. The main goal of a calibration method is to perform a regression of these constants, illustrated by the relation:
$C_{\mathrm{pos},S_{\mathrm{pos}}}\times S_{\mathrm{pos}}\rightarrow E_{\mathrm{ref.}},$ (1)
@@
where “pos” refers to the $x,y$ and $z$ coordinates of the cell, $S_{\mathrm{pos}}$ the signal given by that cell, $C_{\mathrm{pos},S_{\mathrm{pos}}}$ the (discrete) calibration constants for that cell with signal $S_{\mathrm{pos}}$, and $E_{\mathrm{ref.}}$ is the reference energy deposited in that cell (for example, the true/generator energy). The number of $C_{\mathrm{pos},S_{\mathrm{pos}}}$ depends on the binning chosen for the pos and $S$ variables. The goal of calibration is to find the set of calibration constants that minimize the discrepancy between the calibrated energy (left side of Eq. 1) and the reference (right side of Eq. 1). The difficulty in the estimation of the $C_{\mathrm{pos},S_{\mathrm{pos}}}$ parameters is typically due to the following factors:
  1. #1
    The ability to have a direct geometrical correspondence between the cell that one wishes to calibrate and the energy deposit of the whole shower in the calorimeter  [^1] Indeed, it is rare that a particle deposits all its energy in a single cell – this is even more pronounced if the calorimeter is finely segmented and/or inhomogeneous.;
  2. #2
    The knowledge of $E_{\mathrm{ref.}}$ (indeed, the true energy deposit is not often a known quantity), or at least the existence of reasonably close estimator which is independently measured, for example by another sub-detector, such as the tracker in particle physics experiments;
  3. #3
    The statistics available to perform a precise regression of all the $C_{\mathrm{pos},S_{\mathrm{pos}}}$ coefficients;
  4. #4
    Other constraints, for example the potentially limited encoding possibilities for the calibration constants.
@@
Taking CMS electromagnetic calorimeter (ECAL) as an example [1], the calibration constants are derived by either laser monitoring, i.e. using a laser as a standard source of light to induce a response in the ECAL cells (lead-tungsten scintillating crystals), or by physics events such as photons from $\pi^{0}$ or $\eta$ meson decays, or electrons from $W$ or $Z$-boson decays. In these methods, the reference object either corresponds to a single cell by construction (e.g. using a laser as a proxy of the light produced by an energy deposit in a single cell) or often needs to be selected so that it corresponds to a single cell (e.g. selecting narrow reconstructed electrons). Compromises are made in either methods: in the case of laser calibration, there could be mismatches between the proxy signal and the physical signal; and in the case of single cell calibration using electrons, hypotheses have to be made on the shower containment within a single cell. Such compromises are typically challenging, but become likely impossible to reach when high-granularity sampling calorimeters with millions of channels are considered, for instance the High-Granularity Calorimeters (HGCAL) of the CMS experiment [2]. Indeed, it is unimaginable to design a system that would point a laser in every cell of the detector (e.g HGCAL contains $\sim$6 million cells), and the sample of hard electrons which deposit their whole energies in a single cell most likely approaches zero (in the case of HGCAL, multiple-GeV electrons deposit their energies in $\mathcal{O}(100)$ cells).
@@
The Calibr-A-Ton method, described in section 2, aims at overcoming some of those limitations. To evaluate the performance of that method, a high granularity sampling calorimeter design, that consists of a hexagonal electromagnetic section – aimed at the detection of electrons and photons – and two hexagonal hadronic sections placed in succession – aimed at the detection of hadrons – was developed within the GEANT4 software [3]. This simple calorimeter design is described shortly in section 3. In section 4, the performance of Calibr-A-Ton is discussed where known biases in the cell energy response of the calorimeter are introduced, to evaluate the capability of the method to correct for those biases. For illustration, the Calibr-A-Ton procedure is compared to a method where the calibration constants are derived using a lower dimensionality regression. A conclusion is given in section 5.
@@ 2 Description of the Calibr-A-Ton method
The relation 1 can be generalized to all cells belonging to the particle shower:
$\sum_{i=1}^{i=n_{\mathrm{cells}}}C_{\mathrm{pos^{i},S^{i}_{\mathrm{pos}}}}\times S^{i}_{\mathrm{pos}}\rightarrow E_{\mathrm{ref.}},$ (2)
where $i$ are labels for the cells [^2] The number of cells that the method considers as belonging to the particle shower is a tunable parameter: considering a single cell may not capture the whole energy of the shower, while considering too many might lead to integrating more deposits from other particles or more electronics noise. in which a non-vanishing signal, i.e. an energy deposit, is detected. Using this method, the negative effects of factors #1-3 above are drastically reduced. The difficulty from factor #1 is essentially lifted: there is no more need for a direct correspondence between a single cell and the whole energy shower. Furthermore, it is much easier to link the calorimeter signals $S^{i}_{\mathrm{pos}}$ with $E_{\mathrm{ref.}}$, because the reference energy can now be taken as the whole shower energy, and can more easily be independently measured by another sub-detector such as the tracker (factor #2). And finally, since every shower allows to constrain multiple $C_{\mathrm{pos,S_{\mathrm{pos}}}}$ values, less statistics is necessary to obtain an arbitrary precision on the calibration constants (factor #3). Another effect reduces the impact of factor #3 in the new method: there is no need to tightly select showers with given topological properties (for example selecting deposits that occur in a single cell or in low number of cells) – many more showers (in particular high cell multiplicity showers) can be used, thus increasing the statistics.
@@
=^$\mathscr{L}\;\hat{=}\;\sum_{j=0}^{j=N_{\mathrm{showers}}}\frac{\left|\sum_{i}C_{\mathrm{pos^{i,j},S^{i,j}_{pos}}}\times S^{i,j}_{\mathrm{pos}}-E^{j}_{\mathrm{ref.}}\right|}{E^{j}_{\mathrm{ref.}}},$ (3)
@@
The performance of the Calibr-A-Ton method, in which the calibration constants are regressed in bins of the cells energies and 3-dimensional positions, is compared to a standard approach, frequently used in sampling calorimeters, where only a single calibration constant per layer is derived (called per-layer calibration). In that method, the loss function of Eq. 3 simplifies to:
=^$\mathscr{L}_{\ell}\;\hat{=}\;\sum_{j=0}^{j=N_{\mathrm{showers}}}\frac{\left|\sum_{\ell}C_{\ell}\times S^{j}_{\ell}-E^{j}_{\mathrm{ref.}}\right|}{E^{j}_{\mathrm{ref.}}},$ (4)
@@
3 Calorimeter simulation, energy biases and resolution
@@ 3.1 Calorimeter concept
The high-granularity sampling calorimeter design used in this paper, called d2, comes from the OGCID project and is inspired from the CMS HGCAL. As illustrated in Figures 1 and 2 (left), the d2 design features a fully hexagonal geometry, and includes:
  • An electromagnetic section (EE), aimed at the detection of electrons and photons, which consists of 26 successive layers of 320 $\mathrm{\mu m}$ of silicon (the active sensor elements) plus 6.05 $\mathrm{mm}$ of lead (the absorber), totalizing $26$ radiation lengths ($X_{0}$). Each layer features 3571 hexagonal silicon cells [^3] Each EE cell can be inscribed in a 10 mm radius circle., arranged concentrically in 35 rings. The layers are 3 mm distant from one another.
  • Two hadronic sections, one directly behind EE called FH, and one in succession called BH, aimed at the detection of hadrons. The FH (BH) consists of 12 (12) layers of 320 $\mathrm{\mu m}$ of silicon (the active sensor elements) plus 45 $\mathrm{mm}$ (80 mm) of stainless steel (the absorber). Each layer features 1801 hexagonal silicon cells [^4] Each FH/BH cell can be inscribed in a 20 mm radius circle., arranged concentrically in 25 rings. In both FH and BH, the layers are 3 mm distant from one another.
Figure 2 (right) illustrates the interaction of a electron in the EE of the calorimeter, that produces a high cell multiplicity shower.
Figure 1: Geometry of the simple calorimeter used in this paper. It features three hexagonal sections: EE, aimed at the detection of electrons and photons; and FH and BH, aimed at the detection of hadrons.
@@ 3.2 Simulated events
To assess the performance of the Calibr-A-Ton method, $3.571\times 10^{6}$ electrons of known energies, randomly chosen between 10 and 100 GeV, were simulated and shot perpendicularly towards the EE front surface. The energy spectrum of the simulated electrons is shown in Figure 3. To ensure high enough statistics across the whole detector, the simulation is made such that about 10k electron are shot along a trajectory that passes through each cell of the EE front surface. The signals from each silicon cell is taken as the raw energy deposit in the silicon, and is provided by GEANT4. Only cells with an energy deposit greater than $E_{\mathrm{thr}}=14.91$ keV are considered; the energy of the other ones is set to zero. The threshold chosen corresponds approximately to a tenth of the energy deposited by a minimum ionizing particle (MIP) in a 320 $\mathrm{\mu m}$-thick silicon cell. It is applied to reflect the typical dynamic range found in sampling calorimeters readout. Indeed, MIP are often used to assess the absolute energy scale of the detector, and must therefore be typically measured with a signal to noise ratio greater than five [^5] For example, the electronics noise in HGCAL cells corresponds to roughly 20% of a MIP energy deposit..
@@ 3.3 Correction for the absorber
$\mathrm{response}\;\hat{=}\;\frac{\sum_{i}C_{\ell^{i},r^{i},\theta_{k}^{i},S^{i}_{l,r,\theta_{k}}}\times S^{i}_{l,r,\theta_{k}}}{E_{\mathrm{ref.}}}.$ (5)
The distribution of 1/response, when no calibration constants are applied ($C=1$), is shown in Figure 4. For convenience, the absorber correction, which is roughly equal to 75.9, is applied as a multiplicative factor to every cell signal, such that (on average and up to inter-calibration of the cells, fluctuations and noise effects) the energy of one generator particle is equal to the energy deposited in the calorimeter. In that context, after application of the absorber correction, one can expect the calibration constants to be very close to unity in the absence of any significant biases in the cell energy response. In the following, we assume that the absorber correction is always applied to all cells energies, and therefore it is omitted from equations, unless stated otherwise.
@@ 3.4 Binning
The distributions of the cells energies (corrected for the absorber), layer number [^6] For the energies and layer numbers, only cells that satisfy $S_{\mathrm{pos}}>E_{\mathrm{thr}}$ are considered., geometrical ring number, and geometrical $\theta_{k}$ angle are shown in Figure 5. In order to have sufficient statistics for the regression of the cells calibration constants, those are binned along the 3 spatial dimensions ($\ell,r,\theta_{k}$) and along the cell energy ($S$) direction, thus defining the following labels: $\ell_{b},r_{b},\theta_{b},S_{b}$. To avoid any edge effects due to showers depositing a significant amount of their energies outside of the calorimeter volume, only electrons that are generated towards the innermost 26 rings are considered, and the calibration constants for cells between ring numbers 27 and 35 are set to unity and are not optimized. The number of bins along the $\ell,r<27,\theta_{k},S$ directions is 12, 8, 6, 12 respectively. In the case of $r,\theta_{k}$ and $S$, the bins are chosen through a quantile method, such that the cells population in each bin is approximately equal. To account for potential deposits of energy in the hadronic section of the calorimeter, a cell in the hadronic section is considered to belong to a single geometrical layer, ring, and $\theta_{k}$ bin to which the labels $\ell_{b}=27,r_{b}=9,\theta_{b}=1$ are assigned. Its energy is still binned along the energy direction. To reflect the physical variance in the shower development in the direction of the incoming particle (longitudinal direction) as a function of the particle energy, the bins along the layer direction ($\ell$) depend on the energy bin of the cell ($S_{b}$). The bins edges are chosen such that, for each $S_{b}$ bin, the EE cells population in each $\ell_{b}$ bin is equal, as detailed in Eq. 6:
@@
3.5 Energy biases and resolution
3.5.1 Energy biases
The detector described in the previous section does not include any biases in the cell energy response, in the sense that every piece of material reacts the same to a given energy deposit; the geometry is very symmetric; there is no unaccounted dead material that is typically found in physical calorimeters such as cables, cooling devices, electronics components, that can distort the energy response of one cell with respect to the others. For that reason, the detector is essentially free of biases, and calibration constants are expected to be very close to unity.
@@ 3.5.1 Energy biases
3.5.2 Resolution effects
The energy measured by a calorimeter is affected by resolution effects. Those can come from three sources: statistical fluctuations in the number of charges produced by the shower and/or collected by the cell (also known as the stochastic term); noise due to the electronics or other physical processes such as pileup collisions (also known as the noise term); and imperfections in the shower containment, material inhomogeneities, variations in the dimensions in the calorimeter, unaccounted material in front of or within the calorimeter, etc. (collectively known as the constant term). The relative energy resolution $\sigma_{E}/E$ is usually parametrized as:

### v2

@@ Abstract
The energy calibration of calorimeters at collider experiments, such as the ones at the CERN Large Hadron Collider, is crucial for achieving the experiments physics objectives. Standard calibration approaches have limitations that become more pronounced as detector granularity increases. In this paper we propose a novel calibration procedure to simultaneously calibrate individual detector cells belonging to a particle shower by targeting a well-controlled energy reference. The method bypasses some of the difficulties that exist in more standard approaches, and it is implemented using differentiable programming. Simulated energy deposits in the electromagnetic section of a high-granularity calorimeter are used to study the method and demonstrate its performance. It is shown that the method is able to correct for biases in the energy response.
@@ 1 Introduction
The energy calibration of a calorimeter consists in methods that allow to precisely estimate the energy of particles that deposit energy within its volume. An imprecise calorimeter calibration can induce biases and uncertainties on the measurement of particles energies and hence, can significantly reduce the discovery potential of a physics experiment.
@@
Since calorimeters are segmented into detection sub-components, commonly called “cells”, the first step of an energy calibration consists in finding, for each cell, the factor that allows a precise correspondence between the signal given by that cell and the genuine energy deposit that occurred in that cell. These factors are typically called “calibration constants”. The main goal of a calibration method is to perform a regression of these constants, as illustrated by the relation:
$C_{\mathrm{pos},S_{\mathrm{pos}}}\times S_{\mathrm{pos}}\rightarrow E_{\mathrm{ref}},$ (1)
@@
where “pos” refers to the $x,y$ and $z$ coordinates of the cell, $S_{\mathrm{pos}}$ the signal given by that cell, $C_{\mathrm{pos},S_{\mathrm{pos}}}$ the (discrete) calibration constants for that cell with signal $S_{\mathrm{pos}}$, and $E_{\mathrm{ref}}$ is the reference energy deposited in that cell (for example, the true/generator energy). The number of $C_{\mathrm{pos},S_{\mathrm{pos}}}$ depends on the binning chosen for the pos and $S$ variables. The goal of the calibration is to find the set of calibration constants that minimize the discrepancy between the calibrated energy (left side of Eq. 1) and the reference (right side of Eq. 1).
@@
Taking the CMS electromagnetic calorimeter (ECAL) as an example [1], the calibration constants are derived by laser monitoring, i.e. using a laser as a standard source of light to induce a response in individual ECAL cells (lead-tungsten scintillating crystals), and by physics events such as photons from $\pi^{0}$ or $\eta$ meson decays, or electrons from $W$ or $Z$-boson decays. When using a laser as a proxy of the light produced by an energy deposit, the reference object corresponds, by construction, to a single cell. However, when using physics signals, such one-to-one correspondence does not necessarily exist, since particles tend to deposit their energies in multiple calorimeter cells (typically $\mathcal{O}(10)$ cells for the multiple-GeV electrons used in the CMS ECAL calibration). For that reason, the usual calibration procedures focus on calibrating the cell that collects the most energy. Such compromises become extremely challenging when high-granularity sampling calorimeters with millions of channels are considered, for instance the High-Granularity Calorimeter (HGCAL) of the CMS experiment [2]. For example, it is unimaginable to design a system that would point a laser in every cell of the detector (e.g. HGCAL contains $\sim$6 million cells). In addition, in HGCAL, multiple-GeV electrons deposit their energies in $\mathcal{O}(500)$ cells with very large fluctuations in the location and energy distributions and, in that context, focusing the calibration on the cell that collects the most energy is very statistically inefficient: for every particle there is a large amount of information to be learned from the deposits in the $\mathcal{O}(500)$ other cells.
@@
The Calibr-A-Ton method, described in section 2, aims at overcoming some of those limitations. To evaluate the performance of that method, a high granularity sampling calorimeter design, that consists of a hexagonal electromagnetic section – aimed at the detection of electrons and photons – and two hexagonal hadronic sections placed in succession – aimed at the detection of hadrons – was developed within the GEANT4 software [3]. This simple calorimeter design is described shortly in section 3. Section 4 describes the introduction of known biases on the cell energy response and their effects on the energy resolution of the simulated calorimeter. In section 5, the performance of Calibr-A-Ton is discussed in terms of its capability to correct for the known biases introduced. For illustration, the Calibr-A-Ton procedure is compared to a method where the calibration constants are derived using a regression of lower dimensionality. A conclusion is given in section 6.
@@ 2 Description of the Calibr-A-Ton method
Relation 1 can be generalized to all cells belonging to the particle shower:
$\sum_{i=1}^{n_{\mathrm{cells}}}C_{\mathrm{pos^{i},S^{i}_{\mathrm{pos}}}}\times S^{i}_{\mathrm{pos}}\rightarrow E_{\mathrm{ref}},$ (2)
where $i$ are labels for the cells [^1] The number of cells that the method considers as belonging to the particle shower is a tunable parameter: considering a single cell may not capture the whole energy of the shower, while considering too many might lead to integrating more deposits from other particles or more electronics noise. in which a non-vanishing signal, i.e. an energy deposit, is detected. Using this method, every shower allows to constrain multiple $C_{\mathrm{pos,S_{\mathrm{pos}}}}$ values and therefore less statistics is necessary to obtain an arbitrary precision on the calibration constants. In addition, compared to the usual methods, there is a less stringent need to tightly select showers with given topological properties (for example selecting deposits that occur in a low number of cells) – many more showers (in particular high cell multiplicity showers) can be used, further increasing the statistical power of the calibration procedure.
@@
=^$\mathscr{L}\;\hat{=}\;\sum_{j=0}^{N_{\mathrm{showers}}}\frac{\left|\sum_{i}C_{\mathrm{pos^{i,j},S^{i,j}_{pos}}}\times S^{i,j}_{\mathrm{pos}}-E^{j}_{\mathrm{ref}}\right|}{E^{j}_{\mathrm{ref}}},$ (3)
@@
The Calibr-A-Ton method had already been explored in a different context in the Ph.D. theses of Jona Motta [5] and Elena Vernazza [6].
The performance of Calibr-A-Ton, in which the calibration constants are regressed in bins of the cells energies and 3-dimensional positions, is compared to a standard approach, frequently used in sampling calorimeters, where only a single calibration constant per layer is derived (called per-layer calibration). In that method, the loss function of Eq. 3 simplifies to:
=^$\mathscr{L}_{\ell}\;\hat{=}\;\sum_{j=0}^{N_{\mathrm{showers}}}\frac{\left|\sum_{\ell}C_{\ell}\times S^{j}_{\ell}-E^{j}_{\mathrm{ref}}\right|}{E^{j}_{\mathrm{ref}}},$ (4)
@@
3 Calorimeter simulation and event simulation
@@ 3.1 Calorimeter concept
The high-granularity sampling calorimeter design used in this paper, called d2, comes from the OGCID project [7] and is inspired by the CMS HGCAL. As illustrated in Figures 1 and 2 (left), the d2 design features a fully hexagonal geometry, and includes:
  • An electromagnetic section (EE), aimed at the detection of electrons and photons, which consists of 26 successive layers of 320 $\mathrm{\mu m}$ of silicon (the active sensor elements) plus 6.05 $\mathrm{mm}$ of lead (the absorber), totalizing $28$ radiation lengths ($X_{0}$). Each layer features 3571 hexagonal silicon cells arranged concentrically in 35 rings (each EE cell can be inscribed in a 10 mm radius circle). The layers are 3 mm distant from one another and the gaps are filled by air.
  • Two hadronic sections, one directly behind EE called FH, and one in succession called BH, aimed at the detection of hadrons. The FH (BH) consists of 12 (12) layers of 320 $\mathrm{\mu m}$ of silicon (the active sensor elements) plus 45 $\mathrm{mm}$ (80 mm) of stainless steel (the absorber). Each layer features 1801 hexagonal silicon cells, arranged concentrically in 25 rings (each FH/BH cell can be inscribed in a 20 mm radius circle). In both FH and BH, the layers are 3 mm distant from one another and the gaps are filled by air.
Figure 2 (right) illustrates the interaction of an electron in the EE of the calorimeter, that produces a high cell multiplicity shower.
Figure 1: Geometry of the d2 calorimeter used in this paper. It features three hexagonal sections: EE, aimed at the detection of electrons and photons; and FH and BH, aimed at the detection of hadrons.
@@ 3.2 Simulated events
[…]
@@ 3.3 Correction for the absorber
$\mathrm{response}\;\hat{=}\;\frac{\sum_{i}C_{\ell^{i},r^{i},\theta_{k}^{i},S^{i}_{l,r,\theta_{k}}}\times S^{i}_{l,r,\theta_{k}}}{E_{\mathrm{ref}}}.$ (5)
The distribution of 1/response, when no calibration constants are applied ($C=1$), is shown in Figure 4. For convenience, the absorber correction, which is roughly equal to 75.9, is applied as a multiplicative factor to every cell signal, such that (on average and up to inter-calibration of the cells, fluctuations and noise effects) the energy of one generator particle is equal to the energy deposited in the calorimeter. In that context, after application of the absorber correction, one can expect the calibration constants to be very close to unity in the absence of any significant biases in the cell energy response. In the following, we assume that the absorber correction is always applied to all cells energies, and therefore it is omitted from equations.
@@ 3.4 Binning
The distributions of the cells energies (corrected for the absorber), layer numbers (for the energies and layer numbers, only cells that satisfy $S_{\mathrm{pos}}>E_{\mathrm{thr}}$ are considered), geometrical ring numbers, and geometrical $\theta_{k}$ angles are shown in Figure 5. In order to have sufficient statistics for the regression of the cells calibration constants, those are binned along the 3 spatial dimensions ($\ell,r,\theta_{k}$) and along the cell energy ($S$) direction, thus defining the following labels: $\ell_{b},r_{b},\theta_{b},S_{b}$. To avoid any edge effects due to showers depositing a significant amount of their energies outside of the calorimeter volume, only electrons that are generated towards the innermost 26 rings are considered, and the calibration constants for cells between ring numbers 27 and 35 are set to unity and are not optimized. The number of bins along the $\ell,r<27,\theta_{k},S$ directions is 12, 8, 6, 12 respectively. In the case of $r,\theta_{k}$ and $S$, the bins are chosen through a quantile method, such that the cells population in each bin is approximately equal. To account for potential deposits of energy in the hadronic section of the calorimeter, a cell in the hadronic section is considered to belong to a single geometrical layer, ring, and $\theta_{k}$ bin to which the labels $\ell_{b}=27,r_{b}=9,\theta_{b}=1$ are assigned. Its energy is still binned along the energy direction. To reflect the physical variance in the shower development in the direction of the incoming particle (longitudinal direction) as a function of the particle energy, the bins along the layer direction ($\ell$) depend on the energy bin of the cell ($S_{b}$). The bins edges are chosen such that, for each $S_{b}$ bin, the EE cells population in each $\ell_{b}$ bin is equal, as detailed in Eq. 6:
@@
4 Energy biases and resolution
4.1 Energy biases
The detector described in the previous section does not include any biases in the cell energy response, in the sense that every piece of material reacts the same to a given energy deposit; the geometry is very symmetrical; there is no unaccounted dead material that is typically found in physical calorimeters such as cables, cooling devices, electronics components, that can distort the energy response of one cell with respect to the others. For that reason, the detector is essentially free of biases, and calibration constants are expected to be very close to unity.
@@ 4.1 Energy biases
4.2 Resolution effects
The energy measured of a calorimeter is affected by resolution effects. These can come from three sources: statistical fluctuations in the number of charges produced by the shower and/or collected by the cell (also known as the stochastic term); noise due to the electronics or other physical processes such as pileup collisions (also known as the noise term); and imperfections in the shower containment, material inhomogeneities, variations in the dimensions in the calorimeter, unaccounted material in front of or within the calorimeter, etc. (collectively known as the constant term). The relative energy resolution $\sigma_{E}/E$ is usually parametrized as: